# SPAN: Benchmarking and Improving Cross-Calendar Temporal Reasoning of Large Language Models

### v1

@@ Introduction
Recent advances in large language models (LLMs) have led to substantial progress across a range of reasoning tasks (qiao-etal-2023-reasoning; xu2024faithful; li2024understanding; fan-etal-2025-slam). Among these, temporal reasoning, as a fundamental capability for LLMs to understand real-world dynamics, has garnered increasing attention, leading to the development of multiple evaluation benchmarks (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; fatemi2024test; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; ge2025tremu; wei2025time; saxena2025lost). Nevertheless, most existing efforts focus on temporal reasoning within the Gregorian calendar, neglecting the exploration of LLMs’ capability to reason across calendar systems. In practice, numerous countries and regions employ an alternative calendar system alongside the widely-used Gregorian calendar system (taqizadeh1939various; graumann2015problem). For instance, both Saudi Arabia and China use the Gregorian calendar for official purposes, but Saudi Arabia follows the Islamic calendar for religious observances while China relies on the Chinese lunar calendar for traditional cultural activities. These real-world practices underscore the necessity for reasoning and converting dates across calendar systems, i.e., cross-calendar temporal reasoning, which is crucial for global and multicultural applications of LLMs.
Despite its practical significance, the cross-calendar temporal reasoning capability of LLMs remains unexplored. Critically, there is currently no benchmark for assessing this capability, which hinders further research progress. Moreover, existing temporal reasoning benchmarks are inherently time-invariant, limiting evaluation to a single, fixed reference date when the benchmark was created. This leads to two drawbacks: (1) Temporal Scope Limitation. Given the dynamic nature of time, robust temporal reasoning evaluation should ensure a broad temporal scope by enabling assessments at arbitrary reference dates, which is not yet supported by existing benchmarks. (2) Data Contamination. The ground truth in certain temporal reasoning benchmarks is typically deterministic and publicly available through platforms such as Hugging Face and GitHub. Such accessibility may lead to evaluation data being absorbed into training corpora, potentially biasing evaluations (sainz-etal-2023-nlp; deng-etal-2024-investigating; balloccu2024leak).
@@
Temporal reasoning, the capability to perceive and understand the world’s dynamics, is essential for advancing LLMs toward artificial general intelligence (AGI). Early benchmarks such as TimeQA (chen2021datasetansweringtimesensitivequestions) and TimeDial (qin-etal-2021-timedial) concentrate on time-sensitive question answering and temporal commonsense in dialogue contexts. Recently, a new wave of benchmarks has emerged, targeting more fine-grained, complex temporal reasoning scenarios (chen2021datasetansweringtimesensitivequestions; liska2022streamingqa; kasai2023realtime; tan2023benchmarkingimprovingtemporalreasoning; wei-etal-2023-menatqa; tan-etal-2023-towards; wang2024trambenchmarkingtemporalreasoning; chu-etal-2024-timebench; wang-zhao-2024-tram; fatemi2024test; ge2025tremu; wei2025time; saxena2025lost). For example, TimeBench (chu-etal-2024-timebench) introduces a hierarchical benchmark for evaluating a wide spectrum of temporal reasoning capabilities, ranging from temporal relation classification to duration estimation and multi-hop time-aware question answering, spanning symbolic, commonsense, and event-centric tasks. StreamingQA (liska2022streamingqa) and RealtimeQA (kasai2023realtime) extend the focus to dynamic settings where new information arrives continuously, and LLMs are expected to revise their responses over time. Additionally, Lost in Time (saxena2025losttimeclockcalendar) explores how multimodal LLMs process time-related visual inputs, including analogue clocks and yearly calendars. These benchmarks collectively enable systematic evaluation of LLMs’ temporal reasoning across diverse tasks, complexities, and modalities.
@@
In pursuit of human-level temporal reasoning, recent research has proposed various strategies to enhance LLMs (wei-etal-2023-menatqa; tan-etal-2023-towards; yang-etal-2023-upon; jain2023language; sutimo; xiong2024large; yang-etal-2024-enhancing-temporal). For instance, RemeMo (yang-etal-2023-upon) improves LLMs’ temporal understanding during pre-training by organizing events or sentences according to their temporal order, facilitating the model’s capture of complex temporal relationships. TG-LLM (xiong2024large) applies supervised fine-tuning on a synthetic temporal graph dataset, enabling the model to convert text into structured graphs for more effective temporal reasoning. Additionally, TempReason (yang-etal-2024-enhancing-temporal) leverages reinforcement learning to encourage the model to generate temporally coherent predictions, thereby improving the performance across diverse temporal reasoning tasks. These efforts advance the temporal reasoning capabilities of LLMs across various tasks.
@@
In this work, we focus on a practical and representative cross-calendar temporal reasoning scenario, namely, reasoning between the Gregorian calendar and other calendar systems, which is prevalent in many countries and regions (taqizadeh1939various; graumann2015problem). Accordingly, in our formulation, when either the source calendar $c_{s}$} or the target calendar $c_{t}$} is set to the Gregorian calendar, the other is set to a non-Gregorian calendar.
@@
We evaluate six open- and closed-source LLMs. Specifically, the open-source models include Llama-3.3-70B-Instruct (grattafiori2024llama), DeepSeek-V3-1226 (liu2024deepseek), and Qwen-2.5-72B-Instruct (yang2024qwen2), which are accessed via Hugging Face and deployed using the vLLM framework (kwon2023efficient); the closed-source models include GPT-4o (hurst2024gpt), Claude-3.7-Sonnet (anth2025claude), and Gemini-1.5-Pro (team2024gemini), which are accessed through their APIs. All models are evaluated using greedy decoding with a temperature of $0.0$ to ensure reproducibility.
@@ A. Festival List
Calendar
Festival
Gregorian
calendar
- Halloween
- Christmas Day
- New Year’s Day
- Valentine’s Day
- International Women’s Day
- International Workers’ Day
- International Children’s Day
Chinese
lunar calendar
- Chinese New Year
- Lantern Festival
- Dragon Boat Festival
- Chinese Valentine’s Day
- Ghost Festival
- Mid-Autumn Festival
Islamic
calendar
- Hijri New Year
- Isra and Mi’raj
- Eid al-Fitr
- Eid al-Adha
Persian
calendar
- Persian New Year
- Sizdah Be-dar
- Mehregan Festival
- Tirgan Festival
@@ B. Reliability of the GPT-4o Evaluator
Model
Agreement
Accuracy
Spearman Kendall Cohen’s $\bm{\kappa}$
$0.98$
(12321232/$1260$)
$0.95$ $0.95$ $0.95$
GPT-4o
$0.98$
(205205/$210$)
$0.95$ $0.95$ $0.95$
Claude-3.7-
Sonnet
$0.97$
(203203/$210$)
$0.94$ $0.94$ $0.93$
Gemini-1.5-
Pro
$0.97$
(204204/$210$)
$0.88$ $0.88$ $0.87$
Llama-3.3-
70B-Instruct
$0.96$
(201201/$210$)
$0.88$ $0.88$ $0.88$
DeepSeek-V3
$1.00$
(210210/$210$)
$1.00$ $1.00$ $1.00$
Qwen-2.5-
72B-Instruct
$0.99$
(209209/$210$)
$0.99$ $0.99$ $0.99$
@@ C. Statistical Significance Testing
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro
$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3
$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.06$ [$0.05$, $0.07$] $p$ << $0.001$
@@
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.03$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.18$ [$0.17$, $0.18$] $p$ << $0.001$
Gemini-1.5-Pro
$0.11$ [$0.11$, $0.12$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.22$ [$0.21$, $0.22$] $p$ << $0.001$
Deepseek-V3
$0.22$ [$0.21$, $0.23$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.37$ [$0.36$, $0.38$] $p$ << $0.001$
@@
Model
Mean
Difference
95%Confidence
Interval
P-Value
GPT-4o
$0.16$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet
$0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro
$0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct
$0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3
$0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct
$0.06$ [$0.05$, $0.07$] $p$ << $0.001$

### v2

@@ Introduction
Recent advances in large language models (LLMs) have led to substantial progress across a range of reasoning tasks (Qiao et al. 2023; Xu et al. 2024; Li et al. 2024; Fan et al. 2025). Among these, temporal reasoning, as a fundamental capability for LLMs to understand real-world dynamics, has garnered increasing attention, leading to the development of multiple evaluation benchmarks (Chen et al. 2021; Liska et al. 2022; Kasai et al. 2023; Tan et al. 2023b; Wei et al. 2023; Tan et al. 2023a; Fatemi et al. 2024; Wang and Zhao 2024b; Chu et al. 2024; Wang and Zhao 2024a; Ge et al. 2025; Wei et al. 2025; Saxena et al. 2025a). Nevertheless, most existing efforts focus on temporal reasoning within the Gregorian calendar, neglecting the exploration of LLMs’ capability to reason across calendar systems. In practice, numerous countries and regions employ an alternative calendar system alongside the widely-used Gregorian calendar system (Taqizadeh 1939; Graumann 2015). For instance, both Saudi Arabia and China use the Gregorian calendar for official purposes, but Saudi Arabia follows the Islamic calendar for religious observances while China relies on the Chinese lunar calendar for traditional cultural activities. These real-world practices underscore the necessity for reasoning and converting dates across calendar systems, i.e., cross-calendar temporal reasoning, which is crucial for global and multicultural applications of LLMs.
Despite its practical significance, the cross-calendar temporal reasoning capability of LLMs remains unexplored. Critically, there is currently no benchmark for assessing this capability, which hinders further research progress. Moreover, existing temporal reasoning benchmarks are inherently time-invariant, limiting evaluation to a single, fixed reference date when the benchmark was created. This leads to two drawbacks: (1) Temporal Scope Limitation. Given the dynamic nature of time, robust temporal reasoning evaluation should ensure a broad temporal scope by enabling assessments at arbitrary reference dates, which is not yet supported by existing benchmarks. (2) Data Contamination. The ground truth in certain temporal reasoning benchmarks is typically deterministic and publicly available through platforms such as Hugging Face and GitHub. Such accessibility may lead to evaluation data being absorbed into training corpora, potentially biasing evaluations (Sainz et al. 2023; Deng et al. 2024; Balloccu et al. 2024).
@@
Temporal reasoning, the capability to perceive and understand the world’s dynamics, is essential for advancing LLMs toward artificial general intelligence (AGI). Early benchmarks such as TimeQA (Chen et al. 2021) and TimeDial (Qin et al. 2021) concentrate on time-sensitive question answering and temporal commonsense in dialogue contexts. Recently, a new wave of benchmarks has emerged, targeting more fine-grained, complex temporal reasoning scenarios (Chen et al. 2021; Liska et al. 2022; Kasai et al. 2023; Tan et al. 2023b; Wei et al. 2023; Tan et al. 2023a; Wang and Zhao 2024b; Chu et al. 2024; Wang and Zhao 2024a; Fatemi et al. 2024; Ge et al. 2025; Wei et al. 2025; Saxena et al. 2025a). For example, TimeBench (Chu et al. 2024) introduces a hierarchical benchmark for evaluating a wide spectrum of temporal reasoning capabilities, ranging from temporal relation classification to duration estimation and multi-hop time-aware question answering, spanning symbolic, commonsense, and event-centric tasks. StreamingQA (Liska et al. 2022) and RealtimeQA (Kasai et al. 2023) extend the focus to dynamic settings where new information arrives continuously, and LLMs are expected to revise their responses over time. Additionally, Lost in Time (Saxena et al. 2025b) explores how multimodal LLMs process time-related visual inputs, including analogue clocks and yearly calendars. These benchmarks collectively enable systematic evaluation of LLMs’ temporal reasoning across diverse tasks, complexities, and modalities.
@@
In pursuit of human-level temporal reasoning, recent research has proposed various strategies to enhance LLMs (Wei et al. 2023; Tan et al. 2023a; Yang et al. 2023; Jain et al. 2023; Su et al. 2024; Xiong et al. 2024; Yang et al. 2024b). For instance, RemeMo (Yang et al. 2023) improves LLMs’ temporal understanding during pre-training by organizing events or sentences according to their temporal order, facilitating the model’s capture of complex temporal relationships. TG-LLM (Xiong et al. 2024) applies supervised fine-tuning on a synthetic temporal graph dataset, enabling the model to convert text into structured graphs for more effective temporal reasoning. Additionally, TempReason (Yang et al. 2024b) leverages reinforcement learning to encourage the model to generate temporally coherent predictions, thereby improving the performance across diverse temporal reasoning tasks. These efforts advance the temporal reasoning capabilities of LLMs across various tasks.
@@
In this work, we focus on a practical and representative cross-calendar temporal reasoning scenario, namely, reasoning between the Gregorian calendar and other calendar systems, which is prevalent in many countries and regions (Taqizadeh 1939; Graumann 2015). Accordingly, in our formulation, when either the source calendar $c_{s}$} or the target calendar $c_{t}$} is set to the Gregorian calendar, the other is set to a non-Gregorian calendar.
@@
We evaluate six open- and closed-source LLMs. Specifically, the open-source models include Llama-3.3-70B-Instruct (Grattafiori et al. 2024), DeepSeek-V3-1226 (Liu et al. 2024), and Qwen-2.5-72B-Instruct (Yang et al. 2024a), which are accessed via Hugging Face and deployed using the vLLM framework (Kwon et al. 2023); the closed-source models include GPT-4o (Hurst et al. 2024), Claude-3.7-Sonnet (Anthropic 2025), and Gemini-1.5-Pro (Team et al. 2024), which are accessed through their APIs. All models are evaluated using greedy decoding with a temperature of $0.0$ to ensure reproducibility.
@@ A. Festival List
Calendar Festival
Gregorian calendar - Halloween - Christmas Day - New Year’s Day - Valentine’s Day - International Women’s Day - International Workers’ Day - International Children’s Day
Chinese lunar calendar - Chinese New Year - Lantern Festival - Dragon Boat Festival - Chinese Valentine’s Day - Ghost Festival - Mid-Autumn Festival
Islamic calendar - Hijri New Year - Isra and Mi’raj - Eid al-Fitr - Eid al-Adha
Persian calendar - Persian New Year - Sizdah Be-dar - Mehregan Festival - Tirgan Festival
@@ B. Reliability of the GPT-4o Evaluator
Model Agreement Accuracy Spearman Kendall Cohen’s $\bm{\kappa}$
- $0.98$ (12321232/$1260$) $0.95$ $0.95$ $0.95$
GPT-4o $0.98$ (205205/$210$) $0.95$ $0.95$ $0.95$
Claude-3.7- Sonnet $0.97$ (203203/$210$) $0.94$ $0.94$ $0.93$
Gemini-1.5- Pro $0.97$ (204204/$210$) $0.88$ $0.88$ $0.87$
Llama-3.3- 70B-Instruct $0.96$ (201201/$210$) $0.88$ $0.88$ $0.88$
DeepSeek-V3 $1.00$ (210210/$210$) $1.00$ $1.00$ $1.00$
Qwen-2.5- 72B-Instruct $0.99$ (209209/$210$) $0.99$ $0.99$ $0.99$
@@ C. Statistical Significance Testing
Model Mean Difference 95%Confidence Interval P-Value
GPT-4o $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Claude-3.7-Sonnet $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro $0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct $0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3 $0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct $0.06$ [$0.05$, $0.07$] $p$ << $0.001$
@@
Model Mean Difference 95%Confidence Interval P-Value
GPT-4o $0.03$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet $0.18$ [$0.17$, $0.18$] $p$ << $0.001$
Gemini-1.5-Pro $0.11$ [$0.11$, $0.12$] $p$ << $0.001$
Llama-3.3-70B-Instruct $0.22$ [$0.21$, $0.22$] $p$ << $0.001$
Deepseek-V3 $0.22$ [$0.21$, $0.23$] $p$ << $0.001$
Qwen-2.5-72B-Instruct $0.37$ [$0.36$, $0.38$] $p$ << $0.001$
@@
Model Mean Difference 95%Confidence Interval P-Value
GPT-4o $0.16$ [$0.02$, $0.04$] $p$ << $0.001$
Claude-3.7-Sonnet $0.16$ [$0.15$, $0.17$] $p$ << $0.001$
Gemini-1.5-Pro $0.04$ [$0.03$, $0.05$] $p$ << $0.001$
Llama-3.3-70B-Instruct $0.07$ [$0.06$, $0.07$] $p$ << $0.001$
Deepseek-V3 $0.17$ [$0.16$, $0.19$] $p$ << $0.001$
Qwen-2.5-72B-Instruct $0.06$ [$0.05$, $0.07$] $p$ << $0.001$